# 1S-DAug: One-Shot Data Augmentation for Robust Few-Shot Generalization

### v1

@@ B.6 Training-time versus Test-time Augmentation
of size $m_{\mathrm{base}}(M_{\mathrm{tr}}+1)$.
Proposition B.5 (Training-time augmentation bound on base distribution).
Under the assumptions above, there exists $C_{\mathrm{enc}}>0$ (as in Lemma B.2) such that for any $\rho>0$ and $\delta>0$, with probability at least $1-\delta$ over $S_{\mathrm{base}}\sim P_{\mathrm{base}}^{m_{\mathrm{base}}}$, every encoder $\theta$ satisfies
$R_{\mathrm{base}}(\theta)\;\leq\;\widehat{R}_{S_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}},\rho}(\theta)+\frac{2C_{\mathrm{enc}}L_{\mathrm{enc}}}{\rho}\,\frac{r_{0,\mathrm{base}}^{\mathrm{aug}}}{\sqrt{m_{\mathrm{base}}(M_{\mathrm{tr}}+1)}}+\sqrt{\frac{\log(1/\delta)}{2m_{\mathrm{base}}(M_{\mathrm{tr}}+1)}}.$ (26)
@@
Apply Theorem 5.2 to the score class $\mathcal{G}$ on the augmented sample $S_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}}$, with sample size $m_{\mathrm{base}}(M_{\mathrm{tr}}+1)$ and radius parameter $r_{0,\mathrm{base}}^{\mathrm{aug}}$. Lemma B.2 bounds the empirical Rademacher complexity by $C_{\mathrm{enc}}L_{\mathrm{enc}}r_{0,\mathrm{base}}^{\mathrm{aug}}/\sqrt{m_{\mathrm{base}}(M_{\mathrm{tr}}+1)}$, yielding (26). ∎
To relate base and novel risks, assume there exists a discrepancy functional $\mathrm{disc}_{\mathcal{G}}$ such that for all encoders $\theta$,
$R_{\mathrm{novel}}(\theta)\;\leq\;R_{\mathrm{base}}(\theta)+\mathrm{disc}_{\mathcal{G}}\bigl(P_{\mathrm{base}},P_{\mathrm{novel}}\bigr).$ (27)
Combining (27) with Proposition B.5 upper-bounds $R_{\mathrm{novel}}(\theta)$ via a term controlled by training-time augmentation on $P_{\mathrm{base}}$ plus the discrepancy.
@@
$S_{\mathrm{novel}}=\{(x_{i},y_{i})\}_{i=1}^{m_{\mathrm{novel}}}\sim P_{\mathrm{novel}}^{m_{\mathrm{novel}}}.$ (28)
At test time, generate $M$ augmentations per example, form aggregated difference features $\widehat{\Omega}_{\theta}(x_{i})$, and let $\widehat{S}_{\mathrm{novel}}$ denote the sample with these aggregated features. Let $\hat{r}_{\mathrm{novel}}$ be a radius bound for the aggregated novel features, i.e.,
$\|\widehat{\Omega}_{\theta}(x_{i})\|_{2}\leq\hat{r}_{\mathrm{novel}}\qquad\text{for all }i.$
Applying Theorem 5.2 and Lemma B.2 directly to $\widehat{S}_{\mathrm{novel}}$ yields, with probability at least $1-\delta$,
$R_{\mathrm{novel}}(\theta)\;\leq\;\widehat{R}_{\widehat{S}_{\mathrm{novel}},\rho}(\theta)+\frac{2C_{\mathrm{enc}}L_{\mathrm{enc}}}{\rho}\,\frac{\hat{r}_{\mathrm{novel}}}{\sqrt{m_{\mathrm{novel}}}}+\sqrt{\frac{\log(1/\delta)}{2m_{\mathrm{novel}}}}.$ (29)
Proposition B.4 shows that, in an idealized i.i.d. model with $M$ augmentations per example, the maximum radius of novel features decreases after augmentation with probability at least $1/(M+1)$, which in turn tends to reduce $\hat{r}_{\mathrm{novel}}$ relative to the original radius bound. Besides, assuming that the aggregated samples become closer to the typical prototype (Bai et al., 2025), and considering Lemma B.3, we have test-time augmentation shrink the complexity term and reduce the empirical margin risk on the target distribution $P_{\mathrm{novel}}$ itself.
In contrast, training-time augmentation mainly tightens the bound on $R_{\mathrm{base}}(\theta)$; any bound on $R_{\mathrm{novel}}(\theta)$ obtained via (27) still contains the discrepancy $\mathrm{disc}_{\mathcal{G}}(P_{\mathrm{base}},P_{\mathrm{novel}})$, which can be large when base and novel classes differ substantially. This highlights the comparative advantage of test-time augmentation for few-shot generalisation under distribution shift.
@@ E.1 FSL model training (ProtoNet and FEAT)
Image-prompt adapter (optional).

### v3

@@ B.5 Training-time versus Test-time Augmentation
of size $m_{\mathrm{base}}(M_{\mathrm{tr}}+1)$. Let $P_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}}$ denote the induced training-time augmentation distribution on pairwise examples, and define the corresponding pairwise risk by
$R_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}}(\theta):=\mathbb{P}_{(x,y)\sim P_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}}}\bigl(y\,g_{\theta}(x)\leq 0\bigr).$
Assume moreover that the augmented pairs may be treated as i.i.d. draws from $P_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}}$.
Proposition B.6 (Training-time augmentation bound on the induced augmentation distribution).
Under the assumptions above, there exists $C_{\mathrm{enc}}>0$ (as in Lemma B.2) such that for any $\rho>0$ and $\delta>0$, with probability at least $1-\delta$ over the draw of $S_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}}$, every encoder $\theta$ satisfies
$R_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}}(\theta)\;\leq\;\widehat{R}_{S_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}},\rho}(\theta)+\frac{2C_{\mathrm{enc}}B_{X}L_{\mathrm{enc}}}{\rho}\,\frac{r_{0,\mathrm{base}}^{\mathrm{aug}}}{\sqrt{m_{\mathrm{base}}(M_{\mathrm{tr}}+1)}}+\sqrt{\frac{\log(1/\delta)}{2m_{\mathrm{base}}(M_{\mathrm{tr}}+1)}}.$ (45)
@@
Under the i.i.d. assumption stated above, apply Theorem B.1 to the score class $\mathcal{G}$ on the augmented sample $S_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}}$, with sample size $m_{\mathrm{base}}(M_{\mathrm{tr}}+1)$ and radius parameter $r_{0,\mathrm{base}}^{\mathrm{aug}}$. Lemma B.2 bounds the empirical Rademacher complexity by $C_{\mathrm{enc}}B_{X}L_{\mathrm{enc}}r_{0,\mathrm{base}}^{\mathrm{aug}}/\sqrt{m_{\mathrm{base}}(M_{\mathrm{tr}}+1)}$, yielding (45). ∎
To relate training-time augmentation on base classes to novel-class performance, assume there exists a discrepancy functional $\mathrm{disc}_{\mathcal{G}}$ such that for all encoders $\theta$,
$R_{\mathrm{novel}}(\theta)\;\leq\;R_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}}(\theta)+\mathrm{disc}_{\mathcal{G}}\bigl(P_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}},P_{\mathrm{novel}}\bigr).$ (46)
Combining (46) with Proposition B.6 upper-bounds $R_{\mathrm{novel}}(\theta)$ via a term controlled by training-time augmentation on the induced base augmentation distribution plus the discrepancy.
@@
$\mathcal{T}_{\mathrm{novel}}=\{(x_{i},y_{i})\}_{i=1}^{m_{\mathrm{novel}}}\sim P_{\mathrm{novel}}^{m_{\mathrm{novel}}}.$ (47)
At test time, generate $M$ augmentations per example, form aggregated difference features $\bar{\Omega}_{\theta}(x_{i})$, and let $\bar{\mathcal{T}}_{\mathrm{novel}}$ denote the sample evaluated with the aggregated score $\bar{g}_{\theta}(x):=\beta-\|\bar{\Omega}_{\theta}(x)\|_{2}^{2}$. Let $\hat{r}_{\mathrm{novel}}$ be a radius bound for the aggregated novel features, i.e.,
$\|\bar{\Omega}_{\theta}(x_{i})\|_{2}\leq\hat{r}_{\mathrm{novel}}\qquad\text{for all }i.$
Define the corresponding aggregated test-time risk by
$\bar{R}_{\mathrm{novel}}(\theta):=\mathbb{P}_{(x,y)\sim P_{\mathrm{novel}}}\bigl(y\,\bar{g}_{\theta}(x)\leq 0\bigr),$
and the corresponding empirical margin risk by
$\widehat{R}_{\bar{\mathcal{T}}_{\mathrm{novel}},\rho}(\theta):=\frac{1}{m_{\mathrm{novel}}}\sum_{i=1}^{m_{\mathrm{novel}}}\tau_{\rho}\bigl(y_{i}\,\bar{g}_{\theta}(x_{i})\bigr).$
Applying Theorem B.1 to the aggregated score class $\bar{\mathcal{G}}:=\{\bar{g}_{\theta}:\theta\in\Theta\}$ and applying Lemma B.2 with radius parameter $\hat{r}_{\mathrm{novel}}$ yields, with probability at least $1-\delta$,
$\bar{R}_{\mathrm{novel}}(\theta)\;\leq\;\widehat{R}_{\bar{\mathcal{T}}_{\mathrm{novel}},\rho}(\theta)+\frac{2C_{\mathrm{enc}}B_{X}L_{\mathrm{enc}}}{\rho}\,\frac{\hat{r}_{\mathrm{novel}}}{\sqrt{m_{\mathrm{novel}}}}+\sqrt{\frac{\log(1/\delta)}{2m_{\mathrm{novel}}}}.$ (48)
Proposition B.3 heuristically suggests that, the complexity term decreases with probability larger than 0.5. Moreover, under the ensemble effect, test-time augmentation can simultaneously shrink the empirical risk and the complexity term and on the target distribution $P_{\mathrm{novel}}$ itself. In contrast, training-time augmentation mainly tightens the bound on $R_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}}(\theta)$. Any transfer bound on $R_{\mathrm{novel}}(\theta)$ obtained via (46) still contains the discrepancy term $\mathrm{disc}_{\mathcal{G}}\bigl(P_{\mathrm{base}}^{\mathrm{tr\mbox{-}aug}},P_{\mathrm{novel}}\bigr),$ which can remain large when the induced base augmentation distribution and the novel distribution differ substantially. Thus, test-time augmentation can have a comparative advantage for few-shot generalization with the distribution shift.
@@ E.1 FSL model training (ProtoNet and FEAT)
Image-prompt adapter.